(8.62) In 1992, the FAA conducted 20,000 pre-employment drug tests on job applicants who were to be engaged in safety and security-related jobs, and found that 1000 were positive.
(a) Construct a 95 percent confidence interval for the population proportion of positive drug tests.
(b) May you assume that this is a normal distribution? Explain why. (Check the normality of your proportion!)

Answer by stanbon(75887)

In 1992, the FAA conducted 20,000 pre-employment drug tests on job applicants who were to be engaged in safety and security-related jobs, and found that 1000 were positive.
(a) Construct a 95 percent confidence interval for the population proportion
---
p-hat = 1/20 = 0.05
---
ME = 1.96*sqrt[0.05*0.95/20,000) = 0.003
----
95%CI: 0.05-0.003 < p < 0.05+0.003
---
0.047 < p < 0.053
=======================
Cheers,
Stan H.
=============